# JaneEye: A 12-nm 2K-FPS 18.9-J/Frame Event-based Eye Tracking Accelerator

### v1

@@ V Conclusion
This paper presents JaneEye, a breakthrough event-based eye tracking system addressing critical XR requirements. We developed an ultra-lightweight neural network achieving 2.45-pixel accuracy with only 17.6K parameters through our novel ConvJANET architecture. Our co-designed 12-nm ASIC accelerator delivers exceptional performance: 2000 FPS with 0.5 ms latency and 18.9 J/frame energy consumption. Given efficiency demands for wearable integration, our implementation achieves superior EDP while maintaining competitive accuracy. This work establishes new benchmarks for neuromorphic computing in resource-constrained environments, enabling practical deployment of high-performance eye tracking in next-generation immersive technologies.
@@ Acknowledgement
We thank GlobalFoundries for providing us the 12LP-PLUS PDK through the GF12+ University Partnership Programme.

### v2

@@ V Conclusion
This paper presents JaneEye, a breakthrough event-based eye tracking system addressing critical XR requirements. We developed an ultra-lightweight neural network that achieves 2.45-pixel accuracy with only 17.6K parameters, utilizing our novel ConvJANET architecture. Our co-designed 12-nm ASIC accelerator delivers exceptional performance, achieving 2000 FPS with 0.5 ms latency and an energy consumption of 18.9 J/frame. Given efficiency demands for wearable integration, our implementation achieves superior EDP while maintaining competitive accuracy. This work establishes new benchmarks for neuromorphic computing in resource-constrained environments, enabling practical deployment of high-performance eye tracking in next-generation immersive technologies.
@@ Acknowledgement
We thank GlobalFoundries for providing us the 12LP-PLUS PDK through the GF12+ University Partnership Programme. This work was partially supported by the Dutch Research Council (NWO) under the Talent Programme Veni 2023 scheme in Applied and Engineering Sciences (AES), Grant No. 21132 (Energy-Efficient Real-Time Edge Intelligence for Wearable Healthcare Devices).